What is the y-intercept of - 9x + y =33 ?

Found 2 solutions by edjones, algebrahouse.com:
Answer by edjones(8007)

- 9x + y =33
y=9x+33
y=33 the y intercept (when x=0)
.
Ed

Answer by algebrahouse.com(1659)

What is the y-intercept of - 9x + y = 33?

Slope-intercept form is y = mx + b
m is the slope
b is the y-intercept
-9x + y = 33
y = 9x + 33 {added 9x to both sides}

y-intercept is 33
www.algebrahouse.com